# Unlocking the Multi-modal Potential of CLIP for Generalized Category Discovery

### v2

@@ Abstract
Given unlabelled datasets containing both old and new categories, generalized category discovery (GCD) aims to accurately discover new classes while correctly classifying old classes, leveraging the class concepts learned from labeled samples. Current GCD methods only use a single visual modality of information, resulting in poor classification of visually similar classes. As a different modality, text information can provide complementary discriminative information, which motivates us to introduce it into the GCD task. However, the lack of class names for unlabelled data makes it impractical to utilize text information. To tackle this challenging problem, in this paper, we propose a Text Embedding Synthesizer (TES) to generate pseudo text embeddings for unlabelled samples. Specifically, our TES leverages the property that CLIP can generate aligned vision-language features, converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings. Besides, we employ a dual-branch framework, through the joint learning and instance consistency of different modality branches, visual and semantic information mutually enhance each other, promoting the interaction and fusion of visual and text knowledge. Our method unlocks the multi-modal potentials of CLIP and outperforms the baseline methods by a large margin on all GCD benchmarks, achieving new state-of-the-art. The code will be released at https://github.com/enguangW/GET.
Index Terms:
Generalized category discovery, clustering, classification, vision-language pre-training
@@ 1 Introduction
Deep neural networks trained on large amounts of labeled data have shown powerful visual recognition capabilities [1], while this is heartening, the close-set assumption severely hinder the deployment of the model in practical application scenarios. Recently, novel class discovery (NCD) [2] has been proposed to categorize unknown classes of unlabelled data, leveraging the knowledge learned from labeled data. As a realistic extension to NCD, generalized category discovery (GCD) [3] assumes that the unlabelled data comes from both known and unknown classes, rather than just unknown classes as in NCD. The model needs to accurately discover unknown classes while correctly classifying known classes of the unlabelled data, breaking the close-set limitation, making GCD a challenging and meaningful task.
@@
Previous GCD methods [3, 4, 5, 6, 7] utilize a DINO [8] pre-trained ViT as the backbone network to expect good initial discrimination ability of the model, thereby facilitating fine-tuning on the training data. Although promising results have been achieved, the representation of the DINO pre-trained model focuses on specific aspects of the data, resulting in large differences in the clustering performance for different classes in the downstream dataset. To explore the impact of different types of pre-trained models on GCD clustering performance, we conduct an extensive empirical investigation in §$\S$5.5. A key observation is that various pre-trained models perform inferior on distinguishing certain classes, especially those with visually similar appearances, such as the classes in all fine-grained datasets and some super-class subsets of generic datasets. We argue that this is due to the reliance on single visual modality information. Current GCD methods only utilize a single visual modality of information, even visual features obtained by the powerful CLIP [9] visual backbone are still difficult to generalize certain visual concepts, leading to sub-optimal results (see in Fig. 1 (a)). Inspired by the idea that the textual modality can provide complementary discriminative information, we decided to introduce text information into the GCD task to compensate for the insufficient discriminative of visual concepts.
As we all know, thanks to the cross-modal alignment training of very large-scale image-text pairs, CLIP shows strong zero-shot performance, and demonstrates its powerful generalization ability of multi-modal joint embedding. In order to utilize multi-modal joint embedding for GCD tasks, it is a natural idea to conduct joint clustering training with the visual embedding of images and the text embedding of the corresponding class names. However, the lack of class names for unlabelled data in GCD makes it impractical to use the text encoder, thus locking the multi-modal potential of CLIP for GCD.
@@
To tackle this challenging problem, in this paper, we propose a generative-based method to GEnerate pseudo Text embeddings for unlabelled data, dubbed GET. In particular, we introduce a Text Embedding Synthesizer (TES) module based on the vision-language alignment property of CLIP, producing reliable and modality-aligned pseudo text features. As shown in Fig. 1 (b), TES learns a mapping that transforms image embeddings into text embeddings. Specifically, the TES module converts visual embeddings into tokens of the CLIP’s text encoder, eliminating the need for textual input. To ensure that generated pseudo text embeddings reside in the same text space as real text embeddings and maintain consistency, TES distill knowledge from the real text embeddings generated by corresponding class names for labeled data, thus leading to an implicit alignment for unlabelled data. Besides, TES simultaneously aligns the text embedding and the image embedding of the same instance, enforcing the consistency between language and vision while preventing overfitting to known classes. This training approach renders TES equivalent to a special finetuned CLIP text encoder on a specific dataset with only visual input. From another perspective, our TES can be considered as performing an image captioning task [10].
@@
To fully leverage such aligned multi-modal features in the GCD task, we propose a dual-branch multi-modal joint training strategy with a cross-modal instance consistency objective. One branch focuses on visual information, while the other branch supplements it with text information. Through joint learning on the GCD task, visual and semantic information aspects mutually enhance each other. Furthermore, our cross-modal instance consistency objective enforces the instance have the same relationship in both visual and text modalities with anchors constructed by labeled instances, promoting the interaction and alignment of visual and text embedding space. With the supplementation of text embeddings generated by TES and an appropriate dual-branch training strategy, the multi-modal features correct the classification hyperplane, enhancing discriminative ability while reducing bias issues.
@@
  • Through extensive empirical research, we demonstrate that GCD methods are constrained by pre-trained backbone and the representation of a single visual modality makes current GCD methods have inferior clustering performance on many visually similar classes.
  • To tackle the problem that the text encoder can not be used on the unlabelled data, we propose a TES module converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings for each sample. Besides, through the cross-modal instance consistency objective in our dual-branch framework, information of different modalities mutually enhances each other, producing more discriminative classification prototypes.
  • Our method achieves state-of-the-art results on multiple benchmarks, extending GCD to a multi-modal paradigm. Meanwhile, we evaluated GCD methods on ImageNet-R, demonstrating that GCD based on multi-modal features can generalize across different domains, challenging the traditional assumption that the GCD task is based on the same domain.
@@ 2 Related Works
In this section, we first introduce the task of novel class discovery, followed by a review of the existing research on generalized category discovery. Since our proposed method uses the vision-language pre-trained model, we also briefly survey the development of vision-language pre-training.
2.1 Novel Class Discovery.
@@
Novel class discovery (NCD) can be traced back to KCL [11], where pairwise similarity generated by similarity prediction network guides clustering reconstruction, offering a constructive approach for transfer learning across tasks and domains. Early methods are based on two objectives: pretraining on labeled data and clustering on unlabelled data. RS [12] performs a self-supervised pretraining on both labeled and unlabelled data, alleviating the model’s bias towards known classes. Simultaneously, RS proposes knowledge transfer through rank statistics, which has been widely adopted in subsequent research. [13] proposes a two-branch learning framework with dual ranking statistics, exchanging information through mutual knowledge distillation, which is similar to our approach to some extent. Differently, our two branches focus on semantic and visual information, rather than local and global characteristics in [13]. In order to simplify NCD approaches, UNO [14] recommends optimizing the task with a unified cross-entropy loss using the multi-view SwAV [15] exchange prediction strategy, which sets a new NCD paradigm.
2.2 Generalized Category Discovery.
@@
Recently, generalized category discovery (GCD) [3] extends NCD to a more realistic scenario, where unlabelled data comes from both known and unknown classes. GCD [3] employs a pre-trained vision transformer [16] to provide initial visual representations, fine-tuning the backbone through supervised and self-supervised contrastive learning on the labeled and the entire data. Once the model learns discriminative representations, semi-supervised k-means is used for classification by constraining the correct clustering of labeled samples. As an emerging and realistic topic, GCD gradually gaining attention. PromptCAL [5] propose a two-stage framework to tackle the class collision issue caused by false negatives while enhancing the adaptability of the model on downstream datasets. CLIP-GCD [17] mines text descriptions from a large text corpus to use the text encoder and simply concatenates visual and text features for classification. In contrast, our method focuses on the dataset itself, without introducing additional corpus. SimGCD [4] introduces a parametric classification approach, addressing the computational overhead of GCD clustering while achieving remarkable improvements. Specifically, SimGCD adds a classifier on top of GCD and jointly learns self-distillation and supervised training strategies. In this paper, we try to tackle GCD in a parametric [4] way. Unlike previous works [3, 4, 5, 6, 7] that solely relies on visual information, we employ CLIP to introduce multi-modal information into the task, and outperforms current uni-visual-modality GCD methods by a large margin.
2.3 Vision-Language Pre-training.
@@
Vision-Language pre-training [18, 19] aims to train a large-scale model on extensive image-text data, which, through fine-tuning, can achieve strong performance on a range of downstream visual-language tasks. Some studies [20, 21, 22, 23, 24] achieve improved performance in various image-language tasks by modeling image-text interactions through a fusion approach. However, the need to encode all image-text pairs in the fusion approach makes the inference speed in image-text retrieval tasks slow. Consequently, some studies [9, 25] propose a separate encoding of images and texts, and project image and text embeddings into a joint embedding space through contrastive learning. CLIP [9] uses contrastive training on large-scale image-text pairs, minimizing the distance between corresponding images and texts while simultaneously maximizing the distance between non-corresponding pairs. The strong generalization capabilities and multi-modal properties of CLIP prompt us to introduce it to the GCD Task.
@@ 3.2 Parametric GCD method (SimGCD)
$\mathcal{L}_{scon}=-\frac{1}{|B_{l}|}\sum_{i\in B_{l}}\frac{1}{|\mathcal{N}_{i}|}\sum_{q\in\mathcal{N}_{i}}\log\frac{\exp\left(\frac{{\boldsymbol{h}_{i}}^{\top}{\boldsymbol{h}_{q}}^{\prime}}{{\tau_{c}}}\right)}{\sum_{n\in B_{l}}^{n\neq i}\exp\left(\frac{{\boldsymbol{h}_{i}}^{\top}{\boldsymbol{h}_{n}}^{\prime}}{\tau_{c}}\right)}\,,$ (1)
@@
where $\mathcal{N}_{i}$ denotes the indices of other images with the same semantic label as $\boldsymbol{x}_{i}$ in a batch, $B_{l}$ corresponds to the labeled subset of the mini-batch $B$ and $\tau_{c}$ is a temperature value. For visual embeddings $\boldsymbol{z}_{i}$ and $\boldsymbol{z}_{i}^{\prime}$ of two views $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{i}^{\prime}$ generated by the image encoder, an MLP layer $g(\cdot)$ is used to map $\boldsymbol{z}_{i}$ and $\boldsymbol{z}_{i}^{\prime}$ to high-dimensional embeddings $\boldsymbol{h}_{i}=g(\boldsymbol{z}_{i})$ and $\boldsymbol{h}_{i}^{\prime}=g(\boldsymbol{z}_{i}^{\prime})$. For parametric classification, all labeled data are trained by a cross-entropy loss $\mathcal{L}_{\text{cls}}^{s}$ and all training data are trained by a self-distillation loss $\mathcal{L}_{\text{cls}}^{u}$:
@@ 4 Our Method
In this paper, we propose GET, which addresses the GCD task in a multi-modal paradigm. As shown in Fig. 2, our GET contains two stages. In the first stage, we learn a text embedding synthesizer (TES, in §$\S$4.1) to generate pseudo text embeddings for each sample. In the second stage, a dual-branch multi-modal joint training strategy with cross-modal instance consistency (in §$\S$4.2) is introduced to fully leverage multi-modal features.
@@ 4.1 Text embedding synthesizer
where $\boldsymbol{\hat{z}}_{i}^{t}$ and $\boldsymbol{z}_{i}^{v}$ are $\ell_{2}$-normalised, and $\tau_{a}$ is a temperature parameter. Thus, the align loss is $\mathcal{L}_{align}=\mathcal{L}_{align}^{v}+\mathcal{L}_{align}^{t}$
@@
The overall objective of our text embedding synthesizer is calculated as:
$\mathcal{L}_{TES}=\mathcal{L}_{align}+\mathcal{L}_{distill}\,.$ (8)
The distill loss is used to guide pseudo text embeddings of the network’s output towards the real semantic corresponding space and adapt the model to the distribution of the dataset, while the align loss prevents overfitting to known classes and enforces the consistency between two modalities.
Moreover, we introduce a multi-view strategy for TES. Specifically, we calculate both the align loss $\mathcal{L}_{align}$ and distill loss $\mathcal{L}_{distill}$ for two different views $\boldsymbol{x}_{i}$ and ${\boldsymbol{x}_{i}}^{\prime}$ of the same image in a mini-batch. This further implicitly enhances the instance discriminative nature [29] of our TES training, allowing different views of the same labeled image to generate identical pseudo text embeddings. The generated pseudo text embeddings $\boldsymbol{\hat{z}}_{i}^{t}$ are then used for joint training in the second stage.
@@ 4.2 Dual-branch multi-modal joint training
Intuitively, the introduction of multi-modal information can have a positive impact on the GCD task. Textual information can serve as an effective complement to visual information, enhancing the model’s discriminative capabilities. However, how to effectively utilize visual and text information in the GCD task and make the most of their respective roles remains challenging. In this paper, we propose a dual-branch architecture as illustrated in Fig. 2 (b), which focuses on semantic and visual information, respectively. We employ the same parametric training strategy (in §$\S$3.2) for each branch to promote that the model has aligned and complementary discriminative capabilities for visual and text features of the same class. Furthermore, we introduce a cross-modal instance consistency loss, which constrains the instance relationships of samples in both visual and text spaces, enabling the two branches to learn from each other. We use the indicator $v$ to represent the visual concept while $t$ for the text concept.
Visual-branch.  The objective of the visual branch contains a representation learning part and a parametric classification part. Given a visual embedding $\boldsymbol{z}_{i}^{v}$ of image $\boldsymbol{x}_{i}$ generated by the image encoder, an MLP layer $g(\cdot)$ is used to map $\boldsymbol{z}_{i}^{v}$ to a high-dimensional embedding $\boldsymbol{h}_{i}^{v}=g(\boldsymbol{z}_{i}^{v})$. Meanwhile, we employ a prototypical classifier $\eta(\cdot)$ to generate classification probability distribution $\boldsymbol{p}_{i}^{v}=\eta(\boldsymbol{z}_{i}^{v})$. Simply replace all the high-dimensional embedding $\boldsymbol{h}$ (the subscript is omitted for brevity) in Eq. (1) and Eq. (2) with its corresponding visual branch version $\boldsymbol{h}^{v}$ can obtain the the supervised contrastive loss $\mathcal{L}_{scon}^{v}$ and the self-supervised contrastive loss $\mathcal{L}_{ucon}^{v}$. The overall representation learning loss is balanced with $\lambda$, written as:
@@
For the parametric classification part, just replace $\boldsymbol{p}_{i}$ and ${\boldsymbol{p}_{i}}^{\prime}$ of Eq. (3) and Eq. (4) with $\boldsymbol{p}_{i}^{v}$ and ${\boldsymbol{p}_{i}^{v}}^{\prime}$ can obtain the cross-entropy loss $\mathcal{L}_{\text{cls-v}}^{s}$ and the self-distillation loss $\mathcal{L}_{\text{cls-v}}^{u}$. Thus, the classification loss is $\mathcal{L}_{\text{cls}}^{v}=(1-\lambda)\mathcal{L}_{\text{cls-v}}^{u}+\lambda\mathcal{L}_{\text{cls-v}}^{s}$.
The overall objective of visual branch is as:
@@
Text-branch.  Our text branch simply adopts the same training strategy as the visual branch. That is, in particular, given a text embedding $\boldsymbol{\hat{z}}_{i}^{t}$ generated by TES, we first input it into a fully connected layer to gain a learnable text embedding $\boldsymbol{\hat{z}}_{i}^{tl}$ while change its dimension. Simply replace $\boldsymbol{h}_{i}^{v}$ in the representation learning objective $\mathcal{L}_{\text{rep}}^{v}$ with $\boldsymbol{h}_{i}^{t}=g(\boldsymbol{\hat{z}}_{i}^{tl})$ and replace ${\boldsymbol{p}_{i}^{v}}$ in classification parts $\mathcal{L}_{\text{cls}}^{v}$ with $\boldsymbol{p}_{i}^{t}=\eta(\boldsymbol{\hat{z}}_{i}^{tl})$ can yield the corresponding text objectives $\mathcal{L}_{\text{rep}}^{t}$ and $\mathcal{L}_{\text{cls}}^{t}$. In other words, changing the visual conception indicator $v$ into text conception indicator $tl$ can get the corresponding objectives for the text branch. Thus, the overall objective of our text branch can be formally written as:
$\mathcal{L}_{\text{db}}^{t}=\mathcal{L}_{\text{rep}}^{t}+\mathcal{L}_{\text{cls}}^{t}\,.$ (11)
To mitigate the bias between old and new classes, we extended the mean-entropy regularization [30] to a multi-modal mean entropy regularization $H_{mm}={H}(\overline{p}_{mm},\overline{p}_{mm})$, here $\overline{p}_{mm}$ can calculate by $\overline{p}_{mm}=\frac{1}{2|B|}\sum_{i\in B}\left(\sigma({\boldsymbol{p}_{i}^{v}},\tau_{s})+\sigma(\boldsymbol{p}_{i}^{t},\tau_{s})\right)$. In this way, the prediction probabilities in different modalities for each prototype are constrained to be the same, preventing trivial solution.
Cross-modal instance consistency objective.  In order to enable the two branches to learn from each other while encouraging agreement between two different modes, we propose a cross-modal instance consistency objective (CICO), shown in Fig. 2 (c). Our CICO has the same form of mutual knowledge distillation as  [13], but we distill the instance consistency between the two branches. For each mini-batch $B$, we choose its labeled subset $B_{l}$ containing $K$ categories as anchor samples, calculate the visual and text prototypes for $K$ categories as visual anchors $\mathcal{P}_{v}$ and text anchors $\mathcal{P}_{t}$, respectively. We define the instance relationships in visual and text branch as:
$\begin{split}s_{i}^{v}=&\sigma({\boldsymbol{z}_{i}^{v}}^{\top}\mathcal{P}_{v})\\ s_{i}^{t}=&\sigma({\boldsymbol{\hat{z}}_{i}^{tl}}^{\top}\mathcal{P}_{t})\,.\end{split}$ (12)

### v3

# GET: Unlocking the Multi-modal Potential of CLIP for Generalized Category Discovery

@@ Abstract
Given unlabelled datasets containing both old and new categories, generalized category discovery (GCD) aims to accurately discover new classes while correctly classifying old classes. Current GCD methods only use a single visual modality of information, resulting in a poor classification of visually similar classes. As a different modality, text information can provide complementary discriminative information, which motivates us to introduce it into the GCD task. However, the lack of class names for unlabelled data makes it impractical to utilize text information. To tackle this challenging problem, in this paper, we propose a Text Embedding Synthesizer (TES) to generate pseudo text embeddings for unlabelled samples. Specifically, our TES leverages the property that CLIP can generate aligned vision-language features, converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings. Besides, we employ a dual-branch framework, through the joint learning and instance consistency of different modality branches, visual and semantic information mutually enhance each other, promoting the interaction and fusion of visual and text knowledge. Our method unlocks the multi-modal potentials of CLIP and outperforms the baseline methods by a large margin on all GCD benchmarks, achieving new state-of-the-art. Our code is available at: https://github.com/enguangW/GET.
@@ 1 Introduction
Deep neural networks trained on large amounts of labeled data have shown powerful visual recognition capabilities [26]. Although this is heartening, the close-set assumption severely hinders the deployment of the model in practical application scenarios. Recently, novel class discovery (NCD) [17] has been proposed to categorize unknown classes of unlabelled data, leveraging knowledge learned from labeled data. As a realistic extension to NCD, generalized category discovery (GCD)  [45] assumes that the unlabelled data come from both known and unknown classes, rather than just unknown classes as in NCD. The model needs to accurately discover unknown classes while correctly classifying known classes of the unlabelled data, breaking the close-set limitation, making GCD a challenging and meaningful task.
@@
Previous GCD methods [45, 51, 55, 39, 58] utilize a DINO [5] pre-trained ViT as the backbone network to expect good initial discrimination ability of the model, thereby facilitating fine-tuning on the training data. Although promising results have been achieved, these representations derived from a single visual backbone often struggle with visually similar categories, such as the classes in all fine-grained datasets and some super-class subsets of generic datasets. As shown in Fig. 1 (a), replacing the backbone of the parametric baseline [51] with the powerful CLIP [40] visual encoder still struggles to generalize certain visual concepts, leading to sub-optimal results. Inspired by the idea that the textual modality can provide complementary discriminative information, we decided to introduce text information into the GCD task to compensate for the insufficient discriminative of visual concepts. However, the lack of class names for unlabelled data in GCD makes it impractical to use the text encoder, thus locking the multi-modal potential of CLIP for GCD.
@@
In order to tackle this challenging problem, in this paper, we propose a generative-based method to GEnerate pseudo Text embeddings for unlabelled data, dubbed GET. In particular, we first introduce a Text Embedding Synthesizer (TES) module based on the vision-language alignment property of CLIP, producing reliable and modality-aligned pseudo text features. As shown in Fig. 1 (b), TES learns a mapping that transforms image embeddings into text embeddings. Specifically, TES converts visual embeddings into tokens for the text encoder, eliminating the need for textual input. To mitigate the gap between generated pseudo-text embeddings and real text embeddings, TES distills knowledge from real text embeddings corresponding to labeled data. Additionally, TES aligns text and image for the same instance, enforcing consistency between language and vision while preventing overfitting to known classes. This training approach renders TES equivalent to a special finetuned text encoder with only visual input. From another perspective, our TES can be considered as performing an image captioning task [35].
@@
To leverage such multi-modal features in the GCD task, we propose a dual-branch multi-modal joint training strategy with a cross-modal instance consistency objective. One branch focuses on visual information, while the other branch supplements it with text information. Through joint learning on the GCD task, visual and semantic information aspects mutually enhance each other. Furthermore, our cross-modal instance consistency objective enforces the instance have the same relationship in both visual and text modalities with anchors constructed by labeled instances, promoting the interaction and alignment of visual and text embedding space. With the supplementation of text embeddings generated by TES and an appropriate dual-branch training strategy, the multi-modal features correct the classification hyperplane, enhancing discriminative ability while reducing bias issues.
@@
  • To tackle the problem that the text encoder can not be used on the unlabelled data, we propose a TES module converting visual embeddings into tokens of the CLIP’s text encoder to generate pseudo text embeddings.
  • Through the proposed cross-modal instance consistency objective in our dual-branch framework, information of different modalities mutually enhances each other, producing more discriminative classification prototypes.
  • Our method achieves state-of-the-art results on multiple benchmarks, providing GCD a multi-modal paradigm.
@@ 2 Related Works
Novel Class Discovery (NCD).  NCD can be traced back to KCL [20], where pairwise similarity generated by a similarity prediction network guides clustering reconstruction, offering a constructive approach for transfer learning across tasks and domains. Early methods are based on two objectives: pretraining on labeled data and clustering on unlabelled data. RS [18] performs a self-supervised pretraining on both labeled and unlabelled data, alleviating the model’s bias towards known classes. Simultaneously, RS proposes knowledge transfer through rank statistics, which has been widely adopted in subsequent research. [56] proposes a two-branch learning framework with dual ranking statistics, exchanging information through mutual knowledge distillation, which is similar to our approach to some extent. Differently, our two branches focus on semantic and visual information rather than local and global characteristics in [56]. In order to simplify NCD approaches, UNO [14] recommends optimizing the task with a unified cross-entropy loss using the multi-view SwAV [4] exchange prediction strategy, which sets a new paradigm.
@@
Generalized Category Discovery (GCD).  Recently, GCD [45] extends NCD to a more realistic scenario, where unlabelled data comes from both known and unknown classes. GCD [45] employs a pre-trained vision transformer [12] to provide initial visual representations, fine-tuning the backbone through supervised and self-supervised contrastive learning on the labeled and the entire data. Once the model learns discriminative representations, semi-supervised k-means are used for classification by constraining the correct clustering of labeled samples. As an emerging and realistic topic, GCD is gradually gaining attention. PromptCAL [55] propose a two-stage framework to tackle the class collision issue caused by false negatives while enhancing the adaptability of the model on downstream datasets. SimGCD [51] introduces a parametric classification approach, addressing the computational overhead of GCD clustering while achieving remarkable improvements. Specifically, SimGCD adds a classifier on top of GCD and jointly learns self-distillation and supervised training strategies. $\mu$GCD [47] examines the taxonomy bias in previous methods by introducing the Clevr4 dataset and employs the mean-teacher technique and a more efficient training strategy thus achieving remarkable performance improvements. CLIP-GCD [37] mines text descriptions from a large text corpus to use the text encoder and simply concatenates visual and text features for classification. In contrast, our method focuses on the dataset itself, without introducing additional corpus. Most recently, TextGCD [59] collects many text tags from multiple benchmarks and leverages LLMs to enhance these tags, constructing a visual lexicon. It then generates textual descriptions for each sample based on the similarity between the visual lexicon and visual feature. Different from these methods, our GET employs CLIP to introduce multi-modal information into the task without relying on any additional databases or large models.
@@
Vision-Language Pre-training.  Vision-Language pre-training [13, 7, 28, 6, 10, 16] aims to train a large-scale model on extensive image-text data, which, through fine-tuning, can achieve strong performance on a range of downstream visual-language tasks. Some studies [30, 9, 32, 43, 27] achieve improved performance in various image-language tasks by modeling image-text interactions through a fusion approach. However, the need to encode all image-text pairs in the fusion approach makes the inference speed in image-text retrieval tasks slow. Consequently, some studies [40, 21] propose a separate encoding of images and texts, and project image and text embeddings into a joint embedding space through contrastive learning. CLIP [40] uses contrastive training on large-scale image-text pairs, minimizing the distance between corresponding images and texts while simultaneously maximizing the distance between non-corresponding pairs. The strong generalization capabilities and multi-modal properties of CLIP prompt us to introduce it to the GCD Task.
@@ 3.2 Parametric GCD method (SimGCD)
$\mathcal{L}_{scon}=-\frac{1}{|B_{l}|}\sum_{i\in B_{l}}\frac{1}{|\mathcal{N}_{i}|}\sum_{q\in\mathcal{N}_{i}}\log\frac{\exp\left(\frac{{\boldsymbol{h}_{i}}^{\top}{\boldsymbol{h}_{q}}^{\prime}}{{\tau_{sc}}}\right)}{\sum_{n\in B_{l}}^{n\neq i}\exp\left(\frac{{\boldsymbol{h}_{i}}^{\top}{\boldsymbol{h}_{n}}^{\prime}}{\tau_{sc}}\right)}\,,$ (1)
@@
where $\mathcal{N}_{i}$ denotes the indices of other images with the same semantic label as $\boldsymbol{x}_{i}$ in a batch, $B_{l}$ corresponds to the labeled subset of the mini-batch $B$, $\tau_{sc}$ and $\tau_{c}$ are temperature values. For visual embeddings $\boldsymbol{z}_{i}$ and $\boldsymbol{z}_{i}^{\prime}$ of two views $\boldsymbol{x}_{i}$ and $\boldsymbol{x}_{i}^{\prime}$ generated by the image encoder, an MLP layer $g(\cdot)$ is used to map $\boldsymbol{z}_{i}$ and $\boldsymbol{z}_{i}^{\prime}$ to high-dimensional embeddings $\boldsymbol{h}_{i}=g(\boldsymbol{z}_{i})$ and $\boldsymbol{h}_{i}^{\prime}=g(\boldsymbol{z}_{i}^{\prime})$. For parametric classification, all labeled data are trained by a cross-entropy loss $\mathcal{L}_{\text{cls}}^{s}$ and all training data are trained by a self-distillation loss $\mathcal{L}_{\text{cls}}^{u}$:
@@ 4 Our Method
In this paper, we propose GET, which addresses the GCD task in a multi-modal paradigm. As shown in Fig. 2, our GET contains two stages. In the first stage, we learn a text embedding synthesizer (TES, in Sec. 4.1) to generate pseudo text embeddings for each sample. In the second stage, a dual-branch multi-modal joint training strategy with cross-modal instance consistency (in Sec. 4.2) is introduced to fully leverage multi-modal features.
@@ 4.1 Text embedding synthesizer
where $\boldsymbol{\hat{z}}_{i}^{t}$ and $\boldsymbol{z}_{i}^{v}$ are $\ell_{2}$-normalised, and $\tau_{a}$ is a temperature parameter. Thus, the align loss is $\mathcal{L}_{align}=\mathcal{L}_{align}^{v}+\mathcal{L}_{align}^{t}$.
@@
The overall objective of our text embedding synthesizer is $\mathcal{L}_{TES}=\mathcal{L}_{align}+\mathcal{L}_{distill}$. The distill loss is used to guide pseudo text embeddings of the network’s output towards the real semantic corresponding space and adapt the model to the distribution of the dataset, while the align loss prevents overfitting to known classes and enforces the consistency between two modalities. Moreover, we introduce a multi-view strategy for TES. Specifically, we calculate both the align loss $\mathcal{L}_{align}$ and distill loss $\mathcal{L}_{distill}$ for two different views $\boldsymbol{x}_{i}$ and ${\boldsymbol{x}_{i}}^{\prime}$ of the same image in a mini-batch. This further implicitly enhances the instance discriminative nature [53] of our TES training, allowing different views of the same labeled image to generate identical pseudo text embeddings. The generated pseudo text embeddings $\boldsymbol{\hat{z}}_{i}^{t}$ are then used for joint training in the second stage.
@@ 4.2 Dual-branch multi-modal joint training
Intuitively, the introduction of multi-modal information can have a positive impact on the GCD task. Textual information can serve as an effective complement to visual information, enhancing the model’s discriminative capabilities. However, how to effectively utilize visual and text information in the GCD task and make the most of their respective roles remains challenging. In this paper, we propose a dual-branch architecture as illustrated in Fig. 2 (b), which focuses on semantic and visual information, respectively. We employ the same parametric training strategy (in Sec. 3.2) for each branch to promote that the model has aligned and complementary discriminative capabilities for visual and text features of the same class. Furthermore, we introduce a cross-modal instance consistency loss, which constrains the instance relationships of samples in both visual and text spaces, enabling the two branches to learn from each other. We use the indicator $v$ to represent the visual concept while $t$ for the text concept.
Visual-branch.  The objective of the visual branch contains a representation learning part and a parametric classification part. Given a visual embedding $\boldsymbol{z}_{i}^{v}$ of image $\boldsymbol{x}_{i}$ generated by the image encoder, an MLP layer $g(\cdot)$ is used to map $\boldsymbol{z}_{i}^{v}$ to a high-dimensional embedding $\boldsymbol{h}_{i}^{v}=g(\boldsymbol{z}_{i}^{v})$. Meanwhile, we employ a prototypical classifier $\eta(\cdot)$ to generate classification probability distribution $\boldsymbol{p}_{i}^{v}=\eta(\boldsymbol{z}_{i}^{v})$. Simply replace all the high-dimensional embedding $\boldsymbol{h}$ (the subscript is omitted for brevity) in Eq. 1 and Eq. 2 with its corresponding visual branch version $\boldsymbol{h}^{v}$ can obtain the the supervised contrastive loss $\mathcal{L}_{scon}^{v}$ and the self-supervised contrastive loss $\mathcal{L}_{ucon}^{v}$. The overall representation learning loss is balanced with $\lambda$, written as:
@@
For the parametric classification part, just replace $\boldsymbol{p}_{i}$ and ${\boldsymbol{p}_{i}}^{\prime}$ of Eq. 3 and Eq. 4 with $\boldsymbol{p}_{i}^{v}$ and ${\boldsymbol{p}_{i}^{v}}^{\prime}$ can obtain the cross-entropy loss $\mathcal{L}_{\text{cls-v}}^{s}$ and the self-distillation loss $\mathcal{L}_{\text{cls-v}}^{u}$. Thus, the classification loss is $\mathcal{L}_{\text{cls}}^{v}=(1-\lambda)\mathcal{L}_{\text{cls-v}}^{u}+\lambda\mathcal{L}_{\text{cls-v}}^{s}$.
The overall objective of the visual branch is as follows:
@@
Text-branch.  Our text branch simply adopts the same training strategy as the visual branch. That is, in particular, given a text embedding $\boldsymbol{\hat{z}}_{i}^{t}$ generated by TES, we first input it into a fully connected layer to gain a learnable text embedding $\boldsymbol{\hat{z}}_{i}^{tl}$ while change its dimension. Simply replace $\boldsymbol{h}_{i}^{v}$ in the representation learning objective $\mathcal{L}_{\text{rep}}^{v}$ with $\boldsymbol{h}_{i}^{t}=g(\boldsymbol{\hat{z}}_{i}^{tl})$ and replace ${\boldsymbol{p}_{i}^{v}}$ in classification parts $\mathcal{L}_{\text{cls}}^{v}$ with $\boldsymbol{p}_{i}^{t}=\eta(\boldsymbol{\hat{z}}_{i}^{tl})$ can yield the corresponding text objectives $\mathcal{L}_{\text{rep}}^{t}$ and $\mathcal{L}_{\text{cls}}^{t}$. In other words, changing the visual conception indicator $v$ into text conception indicator $tl$ can get the corresponding objectives for the text branch. Thus, the overall objective of our text branch can be formally written as $\mathcal{L}_{\text{db}}^{t}=\mathcal{L}_{\text{rep}}^{t}+\mathcal{L}_{\text{cls}}^{t}$
To mitigate the bias between old and new classes, we extended the mean-entropy regularization [1] to a multi-modal mean entropy regularization $H_{mm}={H}(\overline{p}_{mm},\overline{p}_{mm})$, here $\overline{p}_{mm}$ can calculate by $\overline{p}_{mm}=\frac{1}{2|B|}\sum_{i\in B}\left(\sigma({\boldsymbol{p}_{i}^{v}},\tau_{s})+\sigma(\boldsymbol{p}_{i}^{t},\tau_{s})\right)$. In this way, the prediction probabilities in different modalities for each prototype are constrained to be the same, preventing trivial solutions.
Cross-modal instance consistency objective.  In order to enable the two branches to learn from each other while encouraging agreement between two different modes, we propose a cross-modal instance consistency objective (CICO), shown in Fig. 2 (c). Our CICO has the same form of mutual knowledge distillation as  [56], but we distill the instance consistency between the two branches. For each mini-batch $B$, we choose its labeled subset $B_{l}$ containing $K$ categories as anchor samples, calculate the visual and text prototypes for $K$ categories as visual anchors $\mathcal{P}_{v}$ and text anchors $\mathcal{P}_{t}$, respectively. We define the instance relationships in visual and text branches as:
$\begin{split}s_{i}^{v}=&\sigma({\boldsymbol{z}_{i}^{v}}^{\top}\mathcal{P}_{v})\,,\\ s_{i}^{t}=&\sigma({\boldsymbol{\hat{z}}_{i}^{tl}}^{\top}\mathcal{P}_{t})\,.\end{split}$ (10)